Supplemental Operations Report – March 10th

Date: 3/10/2024
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Hab wall heater, solar observatory, perseverance parking brake, tunnel to science dome

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 73%; Avg 81.5%.
VDC Last 24 hours: Max 59.07V; Min 41.51; Avg 51.29V.
Generator run time: 5340.0. Serviced on 3/10/2024, oil changed and air filter changed

Propane Readings:
Station Tank: 76%
Director Tank: 74%
Intern Tank: 77%
Generator Tank: 59%

Hab Static Tank – 315.5 gallons
GreenHab – 164 gallons
Outpost tank – 400 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Sojourner rover used: No
Hours: 197.2
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance parking brake non-functional

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Refueled 3/7
Crew Car used and why, where: Not used
General notes and comments: N/A

Summary of Internet: All nominal.

EVA suits and radios:
Suits: Batteries in suits 3, 4, and 5 were replaced and are functional now. Batteries in suits 7 and 9 to be replaced
Comms: All nominal

Campus wide inspection, if action taken, what and why: Stakes reset (short). Still need to rework mesh and cut new stakes to appropriate height.

Summary of Hab Operations: Wall heater is not functioning. Diagnostic is pending.
Summary of GreenHab Operations: All nominal
Summary of SciDome Operations: All nominal

Summary of Observatories Operations: Solar observatory telescope diagonal repaired. Shutter mechanism is malfunction. The crew is working on repairs with the help of Peter via email.

Summary of RAM Operations: All nominal
Summary of Outpost Operations: All nominal.
Summary of Health and Safety Issues: All nominal.

Copyright © The Mars Society. All rights reserved. | Main Site